# Smallest Suffixient Sets as a Repetitiveness Measure

### v1

@@ Abstract
Suffixient sets are a novel combinatorial object that capture the essential information of repetitive strings in a way that, provided with a random-access mechanism, supports various forms of pattern matching. In this paper we study the size $\chi$ of the smallest suffixient set as a repetitiveness measure: we place it between known measures and study its sensitivity to various string operations.
Keywords:
Repetitive sequences Text compressibility Burrows-Wheeler Transform.
@@ 1 Introduction
In this paper we better characterize $\chi$ as a repetitiveness measure. First, we study how it behaves when the string undergoes updates, showing in particular that it grows by $O(1)$ when appending or prepending symbols, but that it can grow by $\Omega(\log n)$ upon arbitrary edit operations or rotations, and by $\Omega(\sqrt{n})$ when reversing the string. Second, we show that $\chi=O(r)$ on every string family, where $r$ is the number of equal-letter runs of the BWT of the string. We also show that there are string families where $\chi=o(v)$, where $v$ is the size of the smallest lexicographic parse [27] (an alternative to the size of the Lempel-Ziv parse [20], which behaves similarly). In particular, this holds on the Fibonacci strings, where we fully characterize the only 2 smallest suffixient sets of size 4, and further prove that $\chi\leq\sigma+2$ on all substrings of episturmian words over an alphabet of size $\sigma$. Since $v=O(r)$ on all string families, this settles $\chi$ as a strictly smaller measure than $r$, which is a more natural characterization than in terms of the reverse string. We also show that $\chi$ is incomparable with most “copy-paste” based measures [24], as there are families where it is strictly smaller and others where it is strictly larger than any of those measures.
This result relates to the important question of whether a measure $\mu$ is reachable (i.e., one can represent the string within $O(\mu)$ space), accessible (i.e., one can access any string position from an $O(\mu)$-size representation, in sublinear time), or searchable (i.e., one can search for patterns in sublinear time within space $O(\mu)$). Measure $r$ is, curiously, the only one to date being searchable but unknown to be accessible. Now $\chi$ emerges as a measure smaller than $r$, which can search if provided with a mechanism to efficiently access substrings ($r$ does not need access to support searches). Unlike $r$, $\chi$ is yet unknown to be reachable (as its relation to the smallest known reachable measure, the size $b$ of the smallest bidirectional macro scheme [30], remains unknown). As said, it is known that $\gamma=O(\chi)$, but $\gamma$ is also unknown to be reachable.
@@ 2 Preliminaries
[…]
A string $w[1\mathinner{.\,.}n]$ (or simply $w$ if it is clear from the context) of length $|w|=n$ over the alphabet $\Sigma$ is a sequence $w[1]w[2]\dots w[n]$ of symbols where $w[i]\in\Sigma$ for all $i\in[1,n]$. The empty string string of length 00 is denoted $\epsilon$. We denote by $\Sigma^{*}$ the set of all strings over $\Sigma$. Additionally, we let $\Sigma^{+}=\Sigma^{*}\setminus\{\epsilon\}$ and $\Sigma^{k}=\{w\in\Sigma^{*}\,|\,|w|=k\}$. We denote by $w[i\mathinner{.\,.}j]$ the subsequence $w[i]w[i+1]\dots w[j]$. If $x[1\mathinner{.\,.}n]$ and $y[1\mathinner{.\,.}m]$ are strings, we define the concatenation operation applied on $x$ and $y$, as the string obtained by juxtaposing these two strings, that is, $x\cdot y=x[1]x[2]\cdots x[n]y[1]\cdots y[m]=xy$. A string $x$ is a substring of $w$ if $w=yxz$ for some $y,z\in\Sigma^{*}$. A string $x$ is a prefix of $w$ if $w=xy$ for some $y\in\Sigma^{*}$. Analogously, $x$ is a suffix of $w$ if $w=yx$ for some $y\in\Sigma^{*}$. We say that substrings, prefixes, and suffixes are non-trivial if they are different from $w$ and $\epsilon$. The set of substrings of $w$ is denoted by $\mathcal{F}_{w}$. We also let $\mathcal{F}_{w}(k)=\mathcal{F}_{w}\cap\Sigma^{k}$. The reverse of a finite string $w$ is the string $w^{R}=w[n]\cdot w[n-1]\cdots w[1]$. We denote by $\mathcal{R}(w)$ the multiset of rotations of $w[1\mathinner{.\,.}n]$, that is, $\mathcal{R}(w)=\{w[i+1\mathinner{.\,.}n]w[1\mathinner{.\,.}i]\,|\,i\in[1\mathinner{.\,.}n]\}$. Moreover, we let $\mathcal{R}_{x}(w)$ be the multiset of rotations of $w$ prefixed by the string $x$. The Burrows-Wheeler transform (BWT) of a string $w$, denoted $\mathtt{BWT}(w)$, is the transformation of $w$ obtained by collecting the last symbol of all rotations in $\mathcal{R}(w)$ in lexicographic order. The BWT matrix $B(w)$ of $w$ is the $(n\times n)$-matrix where the $i$-th row is the $i$-th rotation of $w$ in lexicographic order.
A right-infinite string w —we use boldface to emphasize its infinite length— over $\Sigma$ is any infinite sequence $\mathbb{Z}^{+}\rightarrow\Sigma$. The set of all infinite strings over $\Sigma$ is denoted $\Sigma^{\omega}$. A substring of w is the finite string $\textbf{w}[i\mathinner{.\,.}j]$ for any $1\leq i\leq j$. A prefix of w is a finite substring of the form $\textbf{w}[1\mathinner{.\,.}n]$ for some $n\geq 0$. The substring complexity function $P_{\textbf{w}}(k):\mathbb{Z}^{+}\cup\{0\}\rightarrow\mathbb{Z}^{+}$ counts the number of distinct substrings of length $k$ in w, for any $k\in\mathbb{Z}^{+}\cup\{0\}$, that is, $P_{\textbf{w}}(k)=|\mathcal{F}_{\textbf{w}}(k)|$. For a finite string $w[1\mathinner{.\,.}n]$, the domain of $P_{w}$ is restricted to $[0\mathinner{.\,.}n]$.
@@ 2.1 Measures of repetitiveness
Figure 1: Relations between relevant repetitiveness measures and how our results place $\chi$ among them. An arrow $\mu_{1}\to\mu_{2}$ means that $\mu_{1}=O(\mu_{2})$ for all strings and, save for $c\to z$, there is a string family where $\mu_{1}=o(\mu_{2})$. The dotted arrows mark only this last condition, so they are not transitive. Measures in light gray nodes are known to be reachable; those in dark gray are accessible and searchable; and $r$ is hatched because it is searchable but unknown to be accessible.
In this work, we will relate $\chi$, in asymptotic terms, with several well-established measures of repetitiveness [24, 26]: $\delta=\max_{k\in[0\mathinner{.\,.}n]}(\mathcal{F}_{w}(k)/k)$ (a measure of string complexity), $\gamma$ (the smallest string attractor), $b$ (the size of the smallest bidirectional macro scheme), $z$ (the size of a Lempel-Ziv parse), $z_{no}$ (the same without allowing phrases to overlap their sources), $z_{e}$ (the size of the greedy LZ-End parse), $z_{end}$ (the size of the minimal LZ-End parse), $v$ (the size of the smallest lexicographic parse), $r$ (the number of equal-letter runs in the BWT of the string), $g$ (the size of the smallest context-free grammar generating only the string), $g_{rl}$ (the same allowing run-length rules), and $c$ (the size of the smallest collage system generating only the string). Except for $\delta$, $\gamma$ and $r$, these measures are said to be copy-paste because they refer to a way of cutting the sequence into chunks that can be copied from elsewhere in the same sequence. Indeed, $\delta$ and $\gamma$ are lower-bound measures, the former known to be unreachable and the latter unknown to date to be reachable; all the others are. The smallest measures known to be accessible (and searchable) are $z_{end}$ and $g_{rl}$, and $r$ is searchable but unknown to be reachable.
@@ 3 Suffixient Sets and the Measure $\chi$
The set of super-maximal extensions of $w$ is $\mathcal{S}_{r}(w)=\{x\in E_{r}(w)\mid\forall y\in E_{r},y=zx\Rightarrow z=\varepsilon\}$. Moreover, we let $\mathtt{sre}(w)=|\mathcal{S}_{r}(w)|$.
@@
Intuitively, a suffixient set is a collection of positions of $[1\mathinner{.\,.}|w|]$ capturing all the right-extensions appearing in $w$. The smallest suffixient sets, which are suffixient sets of minimum size, have also been characterized in terms of super-maximal right-extensions. The next definition simplifies the original one [4, 5].
@@
Let $w[1\mathinner{.\,.}n]\in\Sigma^{*}$. A suffixient set $S\subseteq[1\mathinner{.\,.}n]$ is a smallest suffixient set for $w$ if there is a bijection $pos:\mathcal{S}_{r}\to S$ such that every $x\in\mathcal{S}_{r}$ is a suffix of $w[1\mathinner{.\,.}pos(x)]$.
In its original formulation, the measure $\chi$ is defined over \$-terminated strings. Here, we define $\chi(w)$ with the \$ being implicit, not forming part of $w$.
@@
Let $w\in\Sigma^{*}$ and assume \$$\text{\tt\$}\not\in\mathcal{F}_{w}$. Then, $\chi(w)=|\mathcal{S}|$, where $\mathcal{S}$ is a smallest suffixient set for $w\text{\tt\$}$.
One can see from the above definitions that $\chi$ is well-defined because $\chi(w)=\mathtt{sre}(w\text{\tt\$})$. We will use this relation to prove results on $\chi$ via $\mathtt{sre}$.
@@ 4 Sensitivity of $\chi$ to String Operations
The sensitivity to string operations has been studied for many repetitiveness measures [1, 9, 10, 14, 15, 23, 28, 29]. It is desirable for a repetitiveness measure to not change much upon small changes in the sequence. Some repetitiveness measures are resistant to edit operations. For instance, $b$, $z$ and $g$ can only increase by a multiplicative constant after an edit operation [1], though they increase only by $O(1)$ when prepending or appending a character. On the other hand, $r$ can increase by a $\Theta(\log n)$ factor when appending a character [15, Prop. 37]. Other results have been obtained concerning more complex string operations, like reversing a string [14], or applying a string morphism [9, 10].
@@
    if $xa$ does not appear in $w$ for any $a\neq c$, then $xc$ induces no new right-extensions in $wc$.
@@
    If for some $a\neq b$, $xa$ and $xb$ were both substrings of $w$, and $c\neq a$ and $c\neq b$, then $xc$ is a new right-extension of $wc$ induced by $xc$.
@@
    If $x$ was always followed by $a\neq c$ in $w$ (hence, not a right-extension), then both $xa$ and $xc$ are new right-extensions of $wc$ induced by $xc$.
Cases 1 and 2 induce at most one new super-maximal right-extension in total for all possible $xc$, namely the longest right-extension in $wc$ that is a suffix of $wc$. For Case 3, consider a fixed $a\in\Sigma$. For all the increasing-length suffixes $x_{1}c,x_{2}c,\dots,x_{t}c$ of $wc$ that became right-extensions together with $x_{1}a,x_{2}a,\dots,x_{t}a$, one can see that the latter form a chain of suffixes of $x_{t}a$. Hence, we only have one possible new super-maximal right-extension ending with $a$, namely $x_{t}a$. There cannot be two of these chains for different symbols: if the suffix $x$ is always followed by $a$, there cannot be a suffix $y$ of $x$ always followed by a different symbol $b$, otherwise, $y$ is followed by $a$ within $xa$, a contradiction.∎
@@
The lower bound follows from Lemma 1. For the upper bound, let $cxa$ be the smallest prefix of $cw$ that was not a right-extension of $w$, but is a right-extension of $cw$ (if it exists). This means that $cxa$ does not appear in $w$ (otherwise, it would be a right-extension of $w)$, so no prefix of $cw$ of length $|cxa|$ or more is right-maximal. Hence, all prefixes of $cw$ shorter than $cxa$ were already right-extensions, and all prefixes longer than $cxa$ cannot be right-extensions. Therefore, $cxa$ together with some $cxb$ appearing in $w$ are the only possible new right-extensions in $cw$ with respect to $w$.∎
By letting $c=\text{\tt\$}$ in Lemma 2, we relate $\chi$ to $\mathtt{sre}$. This makes evident the relation between Combinatorics on words [21] with suffixient sets, via the common notion of right-special factors (what we call here right-maximal substrings).
@@
Let $w\in\Sigma^{*}$ and \$$\text{\tt\$}\not\in\mathcal{F}_{w}$. It holds $\mathtt{sre}(w)+1\leq\chi(w)\leq\mathtt{sre}(w)+2$.
@@
Consider $w=\texttt{abaab}$. It holds $\chi(w)=4$ and $\chi(w\texttt{a})=3$. ∎
@@
Let $w[1\mathinner{.\,.}n]$ be a binary de Bruijn string of order $k$. By definition, $w$ contains every binary string of length $k$ as a substring exactly once. As all the possible pairs of strings $x\texttt{a}$ and $x\texttt{b}$ of length $k$ appear in $w$, it follows that all the strings in $\mathcal{F}_{w}(k)$ are right-extensions. Moreover, each $xc$ with $c\in\{\texttt{a},\texttt{b}\}$ of length $k$ is a super-maximal right-extension: otherwise, there would exist some $d\in\{\texttt{a},\texttt{b}\}$ such that $dxc$ and $dx\overline{c}$ are both substrings of $w$, which raises a contradiction since $dx$ cannot appear twice in $w$. Moreover, there are no right-maximal strings of length $k$ or greater, hence, there are no right-extensions of length greater than $k$. It follows that $\mathtt{sre}(w)=|\mathcal{F}_{w}(k)|=2^{k}=\Omega(n)$. ∎
@@
We prove each claim separately by comparing the supermaximal extensions of $w^{\prime}$ before and after performing the string operation on $w^{\prime}$ that yields $w$, for which $\mathtt{sre}(w)=2^{k}$ by Lemma 5.
@@
For Claim 1, note that $\mathtt{sre}(w^{\prime})$ is the same as $\mathtt{sre}(\texttt{a}^{k}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1})$, as prepending the character a multiple times to this string to obtain $w^{\prime}$ never increases $\mathtt{sre}$; it only updates the supermaximal extension $\texttt{a}^{k}$ to $\texttt{a}^{k+1}$ and $\texttt{a}^{k-1}\texttt{b}$ to $\texttt{a}^{k}\texttt{b}$, and so on. For simplicity, we let $w^{\prime}=\texttt{a}^{k}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1}$. The string $w^{\prime}$ does not contain substrings of length $k$ of the form $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[1\mathinner{.\,.}k-2]$, nor the substring $\texttt{a}^{k-2}\texttt{b}$ b. Note that for each of these substrings $y\in\mathcal{F}_{w}(k)$ with $y\not\in\mathcal{F}_{w^{\prime}}(k)$, the other corresponding right-extension $y^{\prime}$ in $w$ sharing a length $k-1$ prefix with $y$ is not a right-extension in $w^{\prime}$. Moreover, note that all the suffixes of length $k-1$ of these $y$ are not suffixes of one another, nor of the length $k-1$ suffixes of any of the substrings $y^{\prime}$ in $w^{\prime}$. Hence, all $k-1$ length binary strings still appear in $w^{\prime}$ as the suffix of some length $k$ substring that remains a right-extension in $w^{\prime}$, and hence, supermaximal extensions of $w^{\prime}$ have to be of length at least $k$. As each string of length $k$ appearing in $w^{\prime}$ is unique, there are no supermaximal extensions of length greater than $k$. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)$ because we are losing $k-1$ pairs of supermaximal extensions of length $k$ with respect to $w$. It follows that by inserting the b in $w^{\prime}$ to yield $w$, $\mathtt{sre}$ increases by $2(k-1)$.

### v2

@@ Abstract
A suffixient set is a novel combinatorial object that captures the essential information of repetitive strings in a way that, provided with a random access mechanism, supports various forms of pattern matching. In this paper, we study the size $\chi$ of the smallest suffixient set as a repetitiveness measure: we place it between known measures and study its sensitivity to various string operations. As a corollary of our results, we give a simple online algorithm to compute smallest suffixient sets.
@@ 1 Introduction
In this paper we better characterize $\chi$ as a repetitiveness measure. First, we study how it behaves when the string undergoes updates, showing in particular that it grows by $O(1)$ when appending or prepending symbols, but that it can grow additively by $\Omega(\log n)$ upon arbitrary edit operations or rotations, and by $\Omega(\sqrt{n})$ when reversing the string. Second, we show that $\chi=O(r)$ on every string family, where $r$ is the number of equal-letter runs of the BWT of the string. We also show that there are string families where $\chi=o(v)$, where $v$ is the size of the smallest lexicographic parse [28] (an alternative to the size of the Lempel-Ziv parse [20], which behaves similarly). In particular, this holds on the Fibonacci strings, where we fully characterize the only 2 smallest suffixient sets of size 4, and further prove that $\chi\leq\sigma+2$ on all substrings of episturmian words over an alphabet of size $\sigma$. Since $v=O(r)$ on all string families, this settles $\chi$ as a strictly smaller measure than $r$, which is a more natural characterization than in terms of the reverse string. We also show that $\chi$ is incomparable with most “copy-paste” based measures [25], as there are families where it is strictly smaller and others where it is strictly larger than any of those measures.
This result relates to the important question of whether a measure $\mu$ is reachable (i.e., one can represent the string within $O(\mu)$ space), accessible (i.e., one can access any string position from an $O(\mu)$-size representation, in sublinear time), or searchable (i.e., one can search for patterns in sublinear time within space $O(\mu)$). Measure $r$ is, curiously, the only one to date being reachable and searchable, but not known to be accessible. Now $\chi$ emerges as a measure smaller than $r$, which can search if provided with a mechanism to efficiently access substrings ($r$ does not need access to support searches). Unlike $r$, $\chi$ is yet not known to be reachable (as its relation to the smallest known reachable measure, the size $b$ of the smallest bidirectional macro scheme [32], remains unknown). As said, it is known that $\gamma=O(\chi)$, but it is unknown whether $\gamma$ is reachable or not.
Our final contribution is a new, extremely simple, online algorithm to compute smallest suffixient sets (and thus $\chi$). While there already exist efficient algorithms to do this [5], our new algorithm processes the string left to right and, at any point, can exhibit a smallest suffixient set for the string it has just consumed. Just as online suffix tree construction [33], our algorithm uses $O(n)$ space and worst-case time in the transdichotomous RAM model over polynomial-size integer alphabets. The best previous algorithm obtains the same result [5], but it is not online and starts from the suffix array and other components of the suffix tree, whereas ours starts from the text and builds the suffix tree at the same time. All linear-time suffix tree construction algorithms run under the same model of computation.
@@ 2 Preliminaries
An ordered alphabet $\Sigma=\{a_{1},\dots,a_{\sigma}\}$ is a finite set of symbols equipped with a total order << such that $a_{1}<a_{2}<\dots<a_{\sigma}$. When $\sigma=2$, we assume $\Sigma=\{\texttt{a},\texttt{b}\}$ with $\texttt{a}<\texttt{b}$. The special symbol \$, if it appears, is always assumed to be the smallest of the alphabet.
A string $w[1\mathinner{.\,.}n]$ (or simply $w$ if it is clear from the context) of length $|w|=n$ over the alphabet $\Sigma$ is a sequence $w[1]w[2]\cdots w[n]$ of symbols where $w[i]\in\Sigma$ for all $i\in[1,n]$. The empty string of length 0 is denoted $\epsilon$. We denote by $\Sigma^{*}$ the set of all strings over $\Sigma$. Additionally, we let $\Sigma^{+}=\Sigma^{*}\setminus\{\epsilon\}$ and $\Sigma^{k}=\{w\in\Sigma^{*}\,|\,|w|=k\}$. We denote by $w[i\mathinner{.\,.}j]$ the substring $w[i]w[i+1]\cdots w[j]$. If $x=x[1\mathinner{.\,.}n]$ and $y=y[1\mathinner{.\,.}m]$ are strings, we define the concatenation operation applied on $x$ and $y$, as the string obtained by juxtaposing these two strings, that is, $x\cdot y=x[1]x[2]\cdots x[n]y[1]\cdots y[m]=xy$. A string $x$ is a substring of $w$ if $w=yxz$ for some $y,z\in\Sigma^{*}$. A string $x$ is a prefix of $w$ if $w=xy$ for some $y\in\Sigma^{*}$. Analogously, $x$ is a suffix of $w$ if $w=yx$ for some $y\in\Sigma^{*}$. We say that substrings, prefixes, and suffixes are non-trivial if they are different from $w$ and $\epsilon$. The set of substrings of $w$ is denoted by $\mathcal{F}_{w}$. We also let $\mathcal{F}_{w}(k)=\mathcal{F}_{w}\cap\Sigma^{k}$. The reverse of a finite string $w$ is the string $w^{R}=w[n]\cdot w[n-1]\cdots w[1]$. We denote by $\mathcal{R}(w)$ the multiset of rotations of $w[1\mathinner{.\,.}n]$, that is, $\mathcal{R}(w)=\{w[i+1\mathinner{.\,.}n]w[1\mathinner{.\,.}i]\,|\,i\in[1\mathinner{.\,.}n]\}$. The Burrows-Wheeler transform (BWT) of a string $w$, denoted $\mathtt{BWT}(w)$, is the transformation of $w$ obtained by collecting the last symbol of all rotations in $\mathcal{R}(w)$ in lexicographic order. The BWT matrix $B(w)$ of $w$ is the $(n\times n)$-matrix where the $i$-th row is the $i$-th rotation of $w$ in lexicographic order.
A right-infinite string w —we use boldface to emphasize its infinite length— over $\Sigma$ is any infinite sequence $\mathbb{Z}^{+}\rightarrow\Sigma$. The set of all infinite strings over $\Sigma$ is denoted $\Sigma^{\omega}$. A substring of w is the finite string $\textbf{w}[i\mathinner{.\,.}j]$ for any $1\leq i\leq j$. A prefix of w is a finite substring of the form $\textbf{w}[1\mathinner{.\,.}n]$ for some $n\geq 0$. The substring complexity function $P_{\textbf{w}}(k):\mathbb{Z}^{+}\cup\{0\}\rightarrow\mathbb{Z}^{+}$ counts the number of distinct substrings of length $k$ in w, for any $k\in\mathbb{Z}^{+}\cup\{0\}$, that is, $P_{\textbf{w}}(k)=|\mathcal{F}_{\textbf{w}}(k)|$. For a finite string $w[1\mathinner{.\,.}n]$, the domain of $P_{w}$ is restricted to $0\mathinner{.\,.}n]$.
@@ 2.1 Measures of Repetitiveness
Figure 1: Relations between relevant repetitiveness measures and how our results place $\chi$ among them. An arrow $\mu_{1}\to\mu_{2}$ means that $\mu_{1}=O(\mu_{2})$ for all strings and, save for $c\to z$, $z_{no}\to z_{end}$, and $z_{end}\to z_{e}$, there is a string family where $\mu_{1}=o(\mu_{2})$. The dotted arrows mark only this last condition, so they are not transitive. Measures in light gray nodes are known to be reachable; those in dark gray are accessible and searchable; and $r$ is hatched because it is searchable but not known to be accessible.
In this work, we will relate $\chi$, in asymptotic terms, with several well-established measures of repetitiveness [25, 27]: $\delta=\max_{k\in[0\mathinner{.\,.}n]}(\mathcal{F}_{w}(k)/k)$ (a measure of substring complexity), $\gamma$ (the smallest string attractor), $b$ (the size of the smallest bidirectional macro scheme), $z$ (the size of the Lempel-Ziv parse), $z_{no}$ (the same without allowing phrases to overlap their sources), $z_{e}$ (the size of the greedy LZ-End parse), $z_{end}$ (the size of the minimal LZ-End parse), $v$ (the size of the smallest lexicographic parse), $r$ (the number of equal-letter runs in the BWT of the string), $g$ (the size of the smallest context-free grammar generating only the string), $g_{rl}$ (the same allowing run-length rules), and $c$ (the size of the smallest collage system generating only the string). Except for $\delta$, $\gamma$ and $r$, these measures are said to be copy-paste because they refer to a way of cutting the sequence into chunks that can be copied from elsewhere in the same sequence. Indeed, $\delta$ and $\gamma$ are lower-bound measures, the former known to be unreachable and the latter not known to date to be reachable; all the others are. The smallest measures known to be accessible (and searchable) are $z_{end}$ and $g_{rl}$, and $r$ is searchable but not known to be accessible.
@@ 3 Suffixient Sets and the Measure $\chi$
The set of super-maximal extensions of $w$ is $\mathcal{S}_{r}(w)=\{x\in E_{r}(w)\mid\forall y\in E_{r}(w),y=zx\Rightarrow z=\varepsilon\}$. Moreover, we let $\mathtt{sre}(w)=|\mathcal{S}_{r}(w)|$.
@@
Intuitively, a suffixient set is a collection of positions of $1\mathinner{.\,.}|w|]$ capturing all the right-extensions appearing in $w$. The smallest suffixient sets, which are suffixient sets of minimum size, have also been characterized in terms of super-maximal right-extensions. The next definition simplifies the original one [4, 5].
@@
Let $w[1\mathinner{.\,.}n]\in\Sigma^{*}$. A suffixient set $S\subseteq[1\mathinner{.\,.}n]$ is a smallest suffixient set for $w$ if there is a bijection $pos:\mathcal{S}_{r}(w)\to S$ such that every $x\in\mathcal{S}_{r}(w)$ is a suffix of $w[1\mathinner{.\,.}pos(x)]$.
In its original formulation, the measure $\chi$ is defined over \$-terminated strings. Here, we define $\chi(w)$ with the \$ being implicit, not being part of $w$.
@@
Let $w\in\Sigma^{*}$ and assume $\text{\tt\textdollar}\not\in\mathcal{F}_{w}$. Then, $\chi(w)=|\mathcal{S}|$, where $\mathcal{S}$ is a smallest suffixient set for $w\text{\tt\textdollar}$.
[…]
@@ 4 Sensitivity of $\chi$ to String Operations
The sensitivity to string operations has been studied for many repetitiveness measures [1, 9, 10, 14, 15, 24, 29, 30]. It is desirable for a repetitiveness measure to not change much upon small changes in the sequence. Some repetitiveness measures are resistant to edit operations. For instance, $b$, $z$ and $g$ can only increase by a multiplicative constant after an edit operation [1], though they can increase only by a $O(1)$ additive factor when prepending or appending a character. On the other hand, $r$ can increase by a $\Theta(\log n)$ factor when appending a character [15, Prop. 37]. Other results have been obtained concerning more complex string operations, like reversing a string [14], or applying a string morphism [9, 10].
@@
    if $xc$ appears in $w$, or if $xa$ does not appear in $w$ for any $a\neq c$, or both, then $xc$ induces no new right-extensions in $wc$;
@@
    if for some $a\neq b$, $xa$ and $xb$ were both substrings of $w$, and $xc$ was not, then $xc$ is a new right-extension of $wc$;
@@
    if $x$ is always followed by $a\neq c$ in $w$ (hence, $xa$ is not a right-extension of $w$), then both $xa$ and $xc$ are new right-extensions of $wc$.
Cases 1 and 2 induce at most one new super-maximal right-extension in total for all possible $xc$, namely the longest right-extension in $wc$ that is a suffix of $wc$. For Case 3, consider a fixed $a\in\Sigma$. For all the increasing-length suffixes $x_{1}c,x_{2}c,\dots,x_{t}c$ of $wc$ that became right-extensions together with $x_{1}a,x_{2}a,\dots,x_{t}a$, one can see that the latter form a chain of suffixes of $x_{t}a$. Hence, we only have one possible new super-maximal right-extension ending with $a$, namely $x_{t}a$. Observe that the chain of suffixes $x_{1}a,x_{2}a,\dots,x_{t}a$ is unique: if the suffix $x$ is always followed by $a$, any suffix $y$ of $x$ is either right-maximal in $w$ (and $y$ falls within Case 2), or it is always followed by an $a$ (because $x$ is always followed by an $a$), i.e. $y=x_{i}$ for some $i\in[1\mathinner{.\,.}t]$. ∎
Lemma 2 yields a remarkably simple algorithm to compute smallest suffixient sets, which we detail in Section 6.
@@
The lower bound follows from Lemma 1. For the upper bound, let cxacxa be the smallest prefix of $cw$ that is not a right-extension of $w$, but is a right-extension of $cw$ (if it exists). This means that cxacxa does not appear in $w$ (otherwise, it would be a right-extension of $w)$, so no prefix of $cw$ of length $|cxa|$ or more is right-maximal. Hence, all prefixes of $cw$ shorter than cxacxa were already right-extensions, and all prefixes longer than cxacxa cannot be right-extensions. Therefore, cxacxa together with some cxbcxb appearing in $w$, are the only possible new right-extensions in $cw$ with respect to ww.∎
By letting $c=\text{\tt\textdollar}\not\in\mathcal{F}_{w}$ in Lemma 2, we relate $\chi$ to $\mathtt{sre}$ (note that $\chi$ is always at least $\mathtt{sre}+1$ because of the new super-maximal extension ending with \$). This makes clear the relation between Combinatorics on words [21] with suffixient sets, via the common notion of right-special factors (what we call here right-maximal substrings).
@@
Let $w\in\Sigma^{*}$. It holds $\mathtt{sre}(w)+1\leq\chi(w)\leq\mathtt{sre}(w)+2$.
@@
Let $w=\texttt{abaab}$. It holds $\mathcal{S}_{r}(w\text{\tt\textdollar})=\{\texttt{aa},\texttt{ab},\texttt{ab\text{\tt\textdollar}},\texttt{aba}\}$ and $\mathcal{S}_{r}(w\texttt{a}\text{\tt\textdollar})=\mathcal{S}_{r}(\texttt{abaaba\text{\tt\textdollar}})=\{\texttt{ab},\texttt{aba\text{\tt\textdollar}},\texttt{abaa}\}$. Hence, $\chi(w)=4$ and $\chi(w\texttt{a})=3$. ∎
@@
Let $w[1\mathinner{.\,.}n]$ be a binary de Bruijn string of order $k$. By definition, $w$ contains every binary string of length $k$ as a substring exactly once. As all the possible pairs of strings $x\texttt{a}$ and $x\texttt{b}$ of length $k$ appear in $w$, it follows that all the strings in $\mathcal{F}_{w}(k)$ are right-extensions. Moreover, each $x\texttt{a}$ and $x\texttt{b}$ of length $k$ are super-maximal right-extensions: otherwise, there would exist some $c\in\{\texttt{a},\texttt{b}\}$ such that $cx\texttt{a}$ and $cx\texttt{b}$ are both substrings of $w$, which raises a contradiction since the $k$-length string $cx$ cannot appear twice in $w$. Moreover, there are no right-maximal strings of length $k$ or greater; hence, there are no right-extensions of length greater than $k$. It follows that $\mathtt{sre}(w)=|\mathcal{F}_{w}(k)|=2^{k}=\Omega(n)$. ∎
@@
Observe that in each claim, $w$ is obtained after performing a string operation on the corresponding $w^{\prime}$: in Claim 1, $w\in{\tt ins}_{\Sigma}(w^{\prime})$; in Claim 2, $w\in{\tt sub}_{\Sigma}(w^{\prime})$; in Claim 3, $w\in{\tt del}_{\Sigma}(w^{\prime})$; in Claim 4, $w\in\mathcal{R}(w^{\prime})$. We prove each claim separately by comparing the super-maximal extensions of $w^{\prime}$ before and after performing the string operation on $w^{\prime}$ that yields $w$, for which $\mathtt{sre}(w)=2^{k}$ by Lemma 5.
@@
For Claim 1, note that $\mathtt{sre}(w^{\prime})$ is the same as $\mathtt{sre}(\texttt{a}^{k}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1})$, as prepending the character a multiple times to this string to obtain $w^{\prime}$ never increases $\mathtt{sre}$; it only updates the super-maximal extension $\texttt{a}^{k}$ to $\texttt{a}^{k+1}$ and $\texttt{a}^{k-1}\texttt{b}$ to $\texttt{a}^{k}\texttt{b}$, and so on. For simplicity, we let $w^{\prime}=\texttt{a}^{k}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1}$. The string $w^{\prime}$ does not contain substrings of length $k$ of the form $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[1\mathinner{.\,.}k-2]$, nor the substring $\texttt{b}\texttt{a}^{k-2}\texttt{b}$. Note that for each of these substrings $y\in\mathcal{F}_{w}(k)$ with $y\not\in\mathcal{F}_{w^{\prime}}(k)$, the other corresponding right-extension $y^{\prime}$ in $w$ sharing a length k1k-1 prefix with $y$ is not a right-extension in $w^{\prime}$. Moreover, note that all the suffixes of length k1k-1 of these $y$ are not suffixes of one another, nor of the length k1k-1 suffixes of any of the substrings $y^{\prime}$ in $w^{\prime}$. Hence, all k1k-1 length binary strings still appear in $w^{\prime}$ as the suffix of some length $k$ substring that remains a right-extension in $w^{\prime}$, and hence, super-maximal extensions of $w^{\prime}$ have to be of length at least $k$. As each string of length $k$ appearing in $w^{\prime}$ is unique, there are no super-maximal extensions of length greater than $k$. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)$ because we are losing k1k-1 pairs of super-maximal extensions of length $k$ with respect to $w$. It follows that by inserting the b in $w^{\prime}$ to yield $w$, $\mathtt{sre}$ increases by 2k22k-2.